OpenStudy (anonymous):

Use L'Hopital's rule to determine the value of the limit. lim as h -> 0 of 1/h*(integral from 0 to h of arctan(x)/x)

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \frac{ 1 }{ h }\int\limits_{0}^{h}\frac{ \tan^{-1} (x) }{ x }dx\]